From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Feb 28 2003 - 00:15:25 EST
------------------------------------------------------------ Date: Fri Feb 28 00:15:25 2003, Shift: 16:00-24:00 Supervisor: FR, with: Log book#40 pages 40 - 40 Shift Summary: ------------------------------------------------------------ 1) Beam cogged at 18:40. A good store with very low background, still going. At the beginning of the store, ZDC rate was ~8K and ToF rate ~ 27K. 2) Data runs Run 8121 with FS at 8deg., 1/8A and MRS at 40deg., 1050B., Short run with 47K trig.6 events. fb3 was off during this run! Run 8126 with same setting, still running at the end of the shift 3) Problems with fb3, fixed by Flemming (see E-log for details). Run plan for next shift: - Continue Run8126, goal is 300K of trig.6 events. - 200K of trig.6 events with FS at 1/4 field. - 200K of trig.6 events with FS at 1/15 field. - 200K of trig.6 events with FS at 1/15 field, polarity B. ------------------------------------------------------------
This archive was generated by hypermail 2.1.5 : Fri Feb 28 2003 - 00:19:51 EST